Ask
your own question, for FREE!
Mathematics
12 Online
you roll two dice. What is the probability of rolling a 3 on the first die and a 5 on the second one?
Still Need Help?
Join the QuestionCove community and study together with friends!
(1-5/6)(1-5/6)
event A = rolling die one and probability that you donot get 3 is P(A)= 5/6, so the prob of getting 3 is 1-p(A)=1/6 event B = rolling die two and prob that you donot get 5 is P(b)=5/6, so the prob of gtting a 5 is 1-p(b)=1/6 The total probability is P(a)P(b) BECOZ these are INDEPENDENT EVENTS = 1/6*1/6=1/36
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
7 hours ago
0 Replies
0 Medals
3 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal